Please advise the longest point to point measurement inside of a 48" square box and how it was calculated?
Answer by askmemath(368)

Consider a square ABCD.Then the longest point to point measurement would be between the vertices A,C and B,D.The line joining these two points is called the Diagonal.
The diagonal can be found by applying Pythagoras' Theorem to the 2 sides of the square.Since all angles in a square are 90degrees,the diagonal becomes the hypotenuse and the sides of the square will become the sides of the triangle.
In this case D^2 = (48)^2 + (48)^2
D^2 = 2x(48)^2
D = 48 Root(2)